OpenStudy (anonymous):

The equation of the line the point (-1,3) and perpendicular to the line x+2y=-6 is

OpenStudy (anonymous):

Do you know slope intercept form?

OpenStudy (anonymous):

y= mx+b

OpenStudy (anonymous):

You want to find the slop of the line \(x+2y=-6\), so you need to put it in slope intercept form.

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

Do you know how to isolate a variable?

OpenStudy (anonymous):

If you isolate y, then you will be in slope intercept form.

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

So you don't know how to solve for y?

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

Okay if you start with \[ x+2y=-6 \]The first thing you wanna do to isolate \(y\) is to get the other terms on the other side of the equation by subtracting them.. In this case, the other term is \(x\). So to get rid of it, we subtract \(x\) from both sides.\[ x+2y - x = -6 -x \]Now the \(x\) and \(-x\) will cancel out.\[ x-x+2y \rightarrow 2y \]The result is \[ 2y=-6-x \]

OpenStudy (anonymous):

But \(y\) is still not isolated because is has the coefficient \(2\) in front of it. To get rid of that, we have to divide both sides by 2.\[ \frac{2y}{2}= \frac{-6-x}{2} \]This simplifies as follows \[ y = -3-\frac{x}{2} \]

OpenStudy (anonymous):

Are you familiar with this process at all?

OpenStudy (anonymous):

so i would have to subtrat -6 from -x right

OpenStudy (anonymous):

Subtract \(-6\) from \(-x\)? We are solving for \(y\), not for \(x\).

OpenStudy (anonymous):

oh okay no im not fimilar with this process

OpenStudy (anonymous):

y=-3-x/2 how do we solve that

OpenStudy (anonymous):

Anyway, we have the following equation \[ y = -\frac{1}{2}x-3 \]Compare it to \[ y=mx+b \]This means \(m=-1/2\) and \(b=-3\)

OpenStudy (anonymous):

These are the same:\[-\frac{x}{2} = -\frac{1}{2}x\]

OpenStudy (anonymous):

But we're not done yet. We still need to find a perpendicular line.

OpenStudy (anonymous):

so the final answer is x-2y=3

OpenStudy (anonymous):

No

OpenStudy (anonymous):

To find a perpendicular line, we use the other slope. Since the slope of the other line is \(-1/2\), then our line must have a slop of \(2\), because that is the opposite of the reciprocal of \(-1/2\)

OpenStudy (anonymous):

So we know the line we're looking for is of the form \[ y=2x+b \]But we need to find out what \(b\) is.

OpenStudy (anonymous):

It must intersect with the point \((-1,3)\). So we plug it in \[ (3) = 2(-1)+b \]Now we have enough information to solve for \(b\).

OpenStudy (anonymous):

so 2*-2+b